# Data-Driven Prescriptive Analytics Applications: A Comprehensive Survey

### v1

@@ VII Survey statistics
VIII RQ1: Problem types
In order to address RQ1 (cf. Section I), problem types handled by the set of surveyed papers are summarized in Table II. For ease of comprehension, problem types addressed in the set of surveyed papers were grouped under ten categories based on their targeted application domain. An extended discussion on individual use cases found within each application domain can be found in Appendix B in the electronic appendix.
Application domain Papers Problem types
Academia [47, 48, 49, 50, 51, 52, 53] Student guidance [51, 52, 50, 48, 47, 49]
@@ VIII RQ1: Problem types
Advertising and Marketing [54, 55, 56, 57, 58, 59, 60, 61, 62] Targeted advertising [54, 55, 56, 57]
@@
Healthcare [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75] Clinical decision making [63, 64, 65, 66, 67, 68]
@@
Human Resources [76, 77, 78, 79, 80, 81, 82] Recruitment decision making [76, 77, 78]
@@
Infrastructure [83, 84, 85, 86, 87, 88, 89, 90, 91] IT network orchestration [83, 84, 85]
@@
Logistics and Transportation [92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107] Airline operations efficiency [92, 93, 94, 95, 96]
@@
Manufacturing [108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127] Proactive maintenance and error correction [108, 109, 110, 111, 112, 113, 114, 115]
@@
Recreational Activities [128, 129, 130] Social media follower recommendation [128]
@@
Retail and Services [131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145] Inventory management [131, 132, 133, 134, 135, 136]
@@
Social Policy [146, 147, 148, 149, 150] Criminal justice system guidance [146, 147, 148]
@@
TABLE II: Papers and problem types associated with each application domain. ”ESG” abbreviates Environmental, Social, and Governance.
Developing the aforementioned categories was done iteratively during the full text screening phase (cf. Section VI), using previous efforts as the initial expectation (e.g., [14, 9]), yet ultimately ending up with an original taxonomy that would fit the set of surveyed papers - merging, splitting, deleting and inventing categories as necessary. Specifically, the heuristic followed was that a category should contain at least three papers, in order to avoid an overly baroque set of categories, yet at most 20%, or 20 papers, in order for the categories to work as a more useful index within the context of this survey.
As Table II demonstrates, DPSA has been applied to a wide variety of application domains and problem types. To explain some the more specialized terms, Uplift analysis is about designing a treatment, e.g., a promotional campaign, including which customers to target, in order to maximize increases (i.e., ”lifts”) in positive metrics, e.g., profit, due to the treatment [60]. Environmental, Social, and Governance (ESG) pillars refer to a set of responsible organizational practices, which may be subject to legal regulations or be highly prioritized by some investors [149].
The sheer variety of problem types covered by Table II suggests that a broad notion of ”business problem solving” (cf. Definition 1) is reflected by the literature as a whole, concerning organizational value creation [151] for many different kinds of stakeholders, be it in hospitals, factories, universities, social services, grocery stores, or elsewhere. Concrete business concerns naturally vary by domain. Yet one recurring theme is the issue of making the right choice under uncertainty, which indeed motivates the need to combine predictive and prescriptive method elements in DPSA solutions, with such methods being the topic of the following sections.
IX RQ2: Method types
This section addresses part of RQ2 by summarizing the array of individual methods used for DPSA in the set of surveyed papers. Later on, Section X addresses RQ2 in terms of how these methods are combined in different workflow patterns. In both Section IX and X, there is a focus on offering new taxonomies and conceptual models to form a structured understanding of existing DPSA methodologies.
@@ IX RQ2: Method types
Figure 4 illustrates the relative frequency of each method type in the set of surveyed papers. As is apparent from the bar chart, DM/ML and OPT dominate. As we shall see in subsequent analyses, more intricate patterns hide behind this heavily aggregated visual representation however.
@@
With respect to the different method types, Figure 5 and 6 summarize all methods used in the set of surveyed papers for prediction and prescription, respectively (cf. Definition 3), whether they are used alone or in combination with others. The set of methods found under each type in this survey is significantly different from previous surveys within PSA (e.g., [9]), which we attribute to our data-driven focus and a substantial amount of newer papers (cf. Section VII).
The following sections describe each method type, the array of options they provide for prediction and prescription, as well as main trends found in the set of surveyed papers. An extended discussion on each individual method and how it maps to specific applications can be found in Appendix C in the electronic appendix. Throughout, we use the convention of capitalizing method names when they refer to the content of Figure 5 or 6.
@@
IX-A Data Mining and Machine Learning (DM/ML)
This method type covers statistical learning and knowledge discovery from historical datasets [152,  p. 5]. As shown in Figure 5, predictive DM/ML encompasses a large array of methods. Indeed, one might find that this subtree reads like a data mining or ML textbook, with comprehensive coverage of most popular algorithms and approaches.
@@ IX-A Data Mining and Machine Learning (DM/ML)
As shown in Figure 6, prescriptive DM/ML also covers a significant number of approaches. Reinforcement Learning, was used in several papers to solve sequential decision problems (e.g., [55, 68]), for instance, to plan fishing routes based on predicted hauls at each location [122]. Techniques from eXplainable AI (XAI) were also used by some authors to map various measures of feature importance to prescriptions, e.g., to suggest remedial actions students at risk of dropping out [50]. Still other authors find ways to generate prescriptions by providing upstream predictions as input to standard DM/ML methods, e.g., a set of Ensemble Learning models [66, 67], Recommender Systems [52], Graph Neural Networks (GNNs) [142, 129] or Nearest Neighbour Queries [108, 77]. For instance, one paper used a Co-Location Pattern Mining algorithm to first predict viable restaurant locations, which were fed as input to a convolutional GNN making the final suggestion [142]. Prescriptive ML Models, a newer approach on the rise, unify prediction and prescription by training custom ML models, typically with problem-specific loss functions, which can then output prescriptions in a single step, based on input problem parameters (e.g., [121, 79, 80, 134]). For example, in one paper staffing capacity management decisions were provided by a model using a Random Forest Kernel function at its core to weigh the similarity between input points and a dataset of known points [79].
IX-B Mathematical Optimization (OPT)
@@
This general observation is reflected in the breadth of methods used in the set of surveyed papers (cf. Figure 6), which covers a broad array of problem structures and solution approaches, including advanced extensions such as Stochastic (e.g., [72, 95]) and Multi-Objective Programming (e.g., [73, 94]). Still, variants of Linear Problem Models were the most pervasive by far, accounting for 26% of surveyed papers. With only a few exceptions [147, 56, 118], these papers all used linear methods involving integer decision variables, such as Integer (e.g., [87, 146]) and Mixed-Integer Programming (e.g., [85, 104]), with Binary Programming being a special case of the former. A staple within Operations Research [32], linear methods can efficiently solve many types of real-world resource allocation problems, although coping with limited scalability due to NP-hardness and the restrictions of a linear model is a recurrent theme among surveyed papers (e.g., [85, 84, 104, 96]). Heuristic Optimization methods, the runner-up used in 11.5% of papers (e.g., [106, 116, 126]), partially circumvents these issues by easing problem assumptions, at the cost of hard solution guarantees. An objective function requiring black-box ML prediction for evaluation is the typical use case for heuristic methods, e.g., to optimize delivery plans [106] or biodiesel properties [127].
IX-C Probabilistic Modelling (PROB)
This method type covers techniques modelling uncertain causal relationships. While some prescriptive methods already discussed, such as Stochastic Programming and Reinforcement Learning (cf. Section IX-B and IX-A) overlap with PROB conceptually, this category is ultimately more closely associated with a set of predictive methods (cf. Figure 5).
@@ IX-C Probabilistic Modelling (PROB)
IX-D Domain Expertise (EXP)
This method type covers the application of reasoning, knowledge, and heuristics with respect to the problem domain. While it is not unusual to apply a bit of EXP when, e.g., doing feature selection in ML [158], this method type rather refers to the application of EXP as a comprehensive approach to prediction or prescription. As seen in Figure 5 and 6, EXP was used for both aforementioned purposes, and quite commonly so, as is evident from Figure 4.
@@ IX-D Domain Expertise (EXP)
Remaining methods are more application-specific in nature, with one paper using Financial Analysis formulas along with some thresholds for desired yields in order to provide bank loaning strategies [145], and another calculating Proportional Anomaly Adjustments based on detected irregularities and material properties, in order to increase reliability of 3D printing in aerospace manufacturing [125]. Surveyed papers doing uplift analysis for Advertising and Marketing (cf. Section VIII) typically used Qini Curve Analysis, a problem-specific technique analyzing a cumulative curve of predicted uplift per customer to determine which ones to target (e.g., [59, 60, 61]).
IX-E Simulation (SIM)
This method type covers the application of simulation models to capture how systems develop over time. The rarest method type among DPSA applications (cf. Figure 4), papers utilizing SIM all use it for providing prescriptions (cf. Figure 6), each using their own problem-bespoke implementations [110, 71, 74, 135, 54, 149]. Providing prescriptions by evaluating different solution candidates in simulations, i.e., by using Simulation-Based Optimization [159], is the strategy followed by these papers. The aforementioned technique has similar advantages to Heuristic Optimization techniques discussed previously (cf. Section IX-B), since it is able to handle black-box predictive models. In addition, it has a natural way of handling stochastic problem models, in that expected outcomes under complex conditions can be reasoned about by using Monte Carlo approximations [160]. Indeed, all surveyed papers utilizing SIM also use stochastic and/or black-box models as part of their methodology. In one example, the authors simulated different scenarios with respect to a stochastic model, in order to determine the minimum replenishment quantity needed for critical groceries to prevent retailer shortages during COVID-19 panic buying [135].
X RQ2: Method patterns
While Section IX only considered RQ2 in terms of individual method usage, this section considers methodological trends in terms of the combination of methods in current DPSA solutions.
X-A Method type combinations
Figure 7 provides an UpSet plot [161], which uses a matrix layout to summarize how the different method types for prediction and prescription (cf. Section IX) are combined in the set of surveyed papers. Unlike Figure 4, this visual representation makes a distinction in terms of whether the different method types were used for providing predictions or prescriptions in the DPSA workflow (cf. Definition 3). How individual papers were associated with each method combination is described in Appendix D in the electronic appendix.
@@ X-A Method type combinations
X-B Two generic DPSA workflow patterns
@@
The PTP/PWP distinction is deemed useful in an information-theoretical sense, since it forms an approximate 50/50 percentage-wise split of the set of surveyed papers. As we shall see, several concrete workflow patterns and methodological choices are furthermore exclusive to either generic workflow pattern. This bisection is therefore in a sense a ”clean cut” and useful for wrapping our heads around an otherwise untamed territory of methodologies.
As the reader may be aware, the notion of a Predict-Then-Optimize workflow pattern already exists within the methodological PSA literature [22], with the same meaning as PTP for mathematical optimization specifically. PTP is simply a generalization to prescriptive method types beyond OPT, according to the diversity of prescriptive methods found in this survey (cf. Section IX). The PWP pattern has to our knowledge never been explicitly named and subjected to formal analysis - possibly since such workflows form newer and rarer contributions to PSA research.
@@
X-C Predict-Then-Prescribe (PTP) workflows
61 surveyed papers, or 58.7% of all surveyed papers, use a PTP workflow in their solutions. Accounting for 39 of these papers, the most widespread instance of PTP by far is the well-known Predict-Then-Optimize variant [22], in which the role of predictive models, usually based on DM/ML (cf. Figure 7), is to provide inputs to a subsequent prescriptive OPT phase, using Exhaustive Search [137, 89], Dynamic Programming [93, 119], Linear Programming [147, 56, 118], Mixed-Integer Programming [85, 132, 73, 70, 104, 115, 99], Integer Programming [87, 146, 84, 94, 81, 49, 96, 141, 76, 69, 95], Convex Programming [123, 138, 124, 131, 130], Non-Convex Programming [133, 103], or Heuristic Optimization [107, 105, 100] (cf. Section IX-B). In concrete terms, this means that the role of the predictive model is to fill in initial blank coefficients in the optimization model, representing demand forecasts or otherwise, before an optimization algorithm is utilized.

### v2

@@ 7 Survey statistics
8 Problem types
In order to address RQ1 (cf. Section 1), problem types handled by the set of surveyed papers are summarized in Table 2. For ease of comprehension, problem types addressed in the set of surveyed papers were grouped under ten categories based on their targeted application domain. An extended discussion on each individual use case found within each application domain can be found in B.
Application domain Problem types
Academia Student guidance [50, 51, 52, 53, 54, 55]
@@ 8 Problem types
Advertising and Marketing Targeted advertising [57, 58, 59, 60]
@@
Healthcare Clinical decision making [66, 67, 68, 69, 70, 71]
@@
Human Resources Recruitment decision making [79, 80, 81]
@@
Infrastructure IT network orchestration [86, 87, 88]
@@
Logistics and Transportation Airline operations efficiency [95, 96, 97, 98, 99]
@@
Manufacturing Proactive maintenance and error correction [111, 112, 113, 114, 115, 116, 117, 118]
@@
Recreational Activities Social media follower recommendation [131]
@@
Retail and Services Inventory management [134, 135, 136, 137, 138, 139]
@@
Social Policy Criminal justice system guidance [149, 150, 151]
@@
Table 2: Papers and problem types associated with each application domain. ”ESG” abbreviates Environmental, Social, and Governance.
Developing the aforementioned categories was done iteratively during the full text screening phase (cf. Section 6), using previous efforts as the initial expectation (e.g., [14, 9]), yet ultimately ending up with an original taxonomy that would fit the set of surveyed papers - merging, splitting, deleting and inventing categories as necessary. Specifically, the heuristic followed was that a category should contain at least three papers, in order to avoid an overly baroque set of categories, yet at most 20%, or 20 papers, in order for the categories to work as a more useful index within the context of this survey.
As Table 2 demonstrates, DPSA has been applied to a wide variety of application domains and problem types. To explain some of the more specialized terms, Uplift analysis is about designing a treatment, e.g., a promotional campaign, including which customers to target, in order to maximize increases (i.e., ”lifts”) in positive metrics, e.g., profit, due to the treatment [63]. Environmental, Social, and Governance (ESG) pillars refer to a set of responsible organizational practices, which may be subject to legal regulations or be highly prioritized by some investors [152].
As Figure 4 illustrates, the relative activity levels among problem domains have varied over time. Notable trends include the larger number of Healthcare papers published in connection to the COVID-19 pandemic in 2021, along with notable increases in activity for a select few domains in 2023: Manufacturing, Logistics and Transportation, along with Advertising and Marketing. This phenomenon could possibly be related to post-COVID shifts in global supply chains and customer behaviours. In any case, DPSA use cases seem to have been motivated by real-world developments over the years.
Figure 4: Stacked bar chart illustrating the number of papers published within each problem domain over time, as well as the distribution of published papers per year.
The sheer variety of problem types covered by Table 2 suggests that a broad notion of ”business problem solving” (cf. Definition 1) is reflected by the literature as a whole, concerning organizational value creation [154] for many different kinds of stakeholders, be it in hospitals, factories, universities, social services, grocery stores, or elsewhere. Still, due to their common core methodology (cf. Definition 3), patterns in terms of choice of methods and workflows can be identified across use cases and application domains, as Sections 9-11 explore further.
9 Method types
This section addresses part of RQ2 by summarizing the array of individual methods used for DPSA in the set of surveyed papers. Later on, Section 10 addresses RQ2 in terms of how these methods are combined in different workflow patterns. In both Sections 9 and 10, there is a focus on offering new taxonomies and conceptual models to form a structured understanding of existing DPSA methodologies.
@@ 9 Method types
Figure 5 illustrates the relative frequency of each method type in the set of surveyed papers. As is apparent from the bar chart, DM/ML and OPT dominate. As Figure 6 furthermore shows, this trend has only become increasingly pronounced in recent years. As we shall see in subsequent analyses, more intricate patterns hide behind this heavily aggregated visual representation however.
@@
Figure 6: Line chart illustrating the number of papers using each method type published over time, as well as the distribution of published papers per year.
With respect to the different method types, Figures 7 and 8 summarize all methods used in the set of surveyed papers for prediction and prescription, respectively (cf. Definition 3), whether they are used alone or in combination with others. Throughout, we use the convention of capitalizing method names when they refer to the content of Figures 7 or 8.
The following sections describe each method type, the array of options they provide for prediction and prescription, as well as main trends found in the set of surveyed papers. An extended discussion on each individual utilized method and how it maps to specific application scenarios can be found in C. Readers needing detailed information on individual methods beyond what the main text supports are advised to refer to this appendix.
We observe in the extended discussion that methodological choices and considerations vary widely due to the practical constraints of individual applications, even within the same use case and problem domain. For this very reason, we avoid making generic, under-fitted, methodological recommendations. Still, we can derive a few general trends per method type, as presented in the main text.
@@
9.1 Data Mining and Machine Learning (DM/ML)
This method type covers statistical learning and knowledge discovery from historical datasets [155,  p. 5]. As shown in Figure 7, predictive DM/ML encompasses a large array of methods. Indeed, one might find that this subtree reads like a data mining or ML textbook, with comprehensive coverage of most popular algorithms and approaches.
@@ 9.1 Data Mining and Machine Learning (DM/ML)
As shown in Figure 8, prescriptive DM/ML also covers a significant number of approaches. Reinforcement Learning, was used in several papers to solve sequential decision problems (e.g., [58, 71]), for instance, to plan fishing routes based on predicted hauls at each location [125]. Techniques from eXplainable AI (XAI) were also used by some authors to map various measures of feature importance to prescriptions, e.g., to suggest remedial actions to students at risk of dropping out [52]. Still other authors find ways to generate prescriptions by providing upstream predictions as input to standard DM/ML methods, e.g., a set of Ensemble Learning models [69, 70], Recommender Systems [51], Graph Neural Networks (GNNs) [145, 132] or Nearest Neighbour Queries [111, 80]. For instance, one paper used a Co-Location Pattern Mining algorithm to first predict viable restaurant locations, which were fed as input to a convolutional GNN making the final suggestion [145]. Prescriptive ML Models, a newer approach on the rise, unify prediction and prescription by training custom ML models, typically with problem-specific loss functions, which can then output prescriptions in a single step, based on input problem parameters (e.g., [124, 82, 83, 137]). For example, in one paper staffing capacity management decisions were provided by a model using a Random Forest Kernel function at its core to weigh the similarity between input points and a dataset of known points [82].
9.2 Mathematical Optimization (OPT)
@@
The challenge of dealing with the aforementioned approximation trade-off is reflected in the breadth of methods used in the set of surveyed papers (cf. Figure 8), which covers a broad array of problem structures and solution approaches, including advanced extensions such as Stochastic (e.g., [75, 98]) and Multi-Objective Programming (e.g., [76, 97]). Still, variants of Linear Problem Models were the most pervasive by far, accounting for 26% of surveyed papers. With only a few exceptions [150, 59, 121], these papers all used linear methods involving integer decision variables, such as Integer (e.g., [90, 149]) and Mixed-Integer Programming (e.g., [88, 107]), with Binary Programming being a special case of the former. A staple within Operations Research [35], linear methods can efficiently solve many types of real-world resource allocation problems with many variables. However, coping with limited scalability due to NP-hardness and the restrictions of a linear model is a recurrent theme among surveyed papers (e.g., [88, 87, 107, 99]). Heuristic Optimization methods, the runner-up used in 11.5% of papers (e.g., [109, 119, 129]), partially circumvents these issues by easing problem assumptions, at the cost of hard solution guarantees (cf. the approximation trade-off). An objective function requiring black-box ML prediction for evaluation is the typical use case for heuristic methods, e.g., to optimize delivery plans [109] or biodiesel properties [130].
9.3 Probabilistic Modelling (PROB)
This method type covers techniques modelling uncertain causal relationships. While some prescriptive methods already discussed, such as Stochastic Programming and Reinforcement Learning (cf. Sections 9.2 and 9.1) overlap with PROB conceptually, this category is ultimately more closely associated with a set of predictive methods (cf. Figure 7).
@@ 9.3 Probabilistic Modelling (PROB)
9.4 Domain Expertise (EXP)
This method type covers the application of reasoning, knowledge, and heuristics with respect to the problem domain. While it is not unusual to apply a bit of EXP when, e.g., doing feature selection in ML [161], this method type rather refers to the application of EXP as a comprehensive approach to prediction or prescription. As seen in Figures 7 and 8, EXP was used for both aforementioned purposes, and quite commonly so, as is evident from Figure 5.
@@ 9.4 Domain Expertise (EXP)
Remaining methods are more application-specific in nature, with one paper using Financial Analysis formulas along with some thresholds for desired yields in order to provide bank loaning strategies [148], and another calculating Proportional Anomaly Adjustments based on detected irregularities and material properties, in order to increase reliability of 3D printing in aerospace manufacturing [128]. Surveyed papers doing uplift analysis for Advertising and Marketing (cf. Section 8) typically used Qini Curve Analysis, a problem-specific technique analyzing a cumulative curve of predicted uplift per customer to determine which ones to target (e.g., [62, 63, 64]).
9.5 Simulation (SIM)
This method type covers the application of simulation models to capture how systems develop over time. The rarest method type among DPSA applications (cf. Figure 5), papers utilizing SIM all use it for providing prescriptions (cf. Figure 8), each using their own problem-bespoke implementations [113, 74, 77, 138, 57, 152]. Providing prescriptions by evaluating different solution candidates in simulations, i.e., by using Simulation-Based Optimization [162], is the strategy followed by these papers. The aforementioned technique has similar advantages to Heuristic Optimization techniques discussed previously (cf. Section 9.2), since it is able to handle black-box predictive models. In addition, it has a natural way of handling stochastic problem models, in that expected outcomes under complex conditions can be reasoned about by using Monte Carlo approximations [163]. Indeed, all surveyed papers utilizing SIM also use stochastic and/or black-box models as part of their methodology. In one example, the authors simulated different scenarios with respect to a stochastic model, in order to determine the minimum replenishment quantity needed for critical groceries to prevent retailer shortages during COVID-19 panic buying [138].
10 Method patterns
While Section 9 only considered RQ2 in terms of individual method usage, this section considers methodological trends in terms of the combination of methods in current DPSA solutions.
10.1 Method type combinations
Figure 9 provides an UpSet plot [164], which uses a matrix layout to summarize how the different method types for prediction and prescription (cf. Section 9) are combined in the set of surveyed papers. Unlike Figure 5, this visual representation makes a distinction in terms of whether the different method types were used for providing predictions or prescriptions in the DPSA workflow (cf. Definition 3). How individual papers were associated with each method combination is described in D.
@@ 10.1 Method type combinations
Similarly, Figure 10 shows that while combinations of DM/ML and OPT have become increasingly dominant over time among the top five most popular method combinations within DPSA, solutions instead using DM/ML or EXP for prescriptions have also become increasingly popular. This implies that methodological exploration has not yet stagnated within the field, despite what the increasing dominance of the DM/ML and OPT method types in Figure 6 might suggest in a vacuum.
Figure 10: Line chart illustrating the relative popularity of the top five most popular method combination over time in terms of the number of published papers per year.
10.2 Two generic DPSA workflow patterns
@@
The PTP/PWP distinction is deemed useful in an information-theoretical sense, since it forms an approximate 50/50 percentage-wise split of the set of surveyed papers. As illustrated in Figure 12, the relative popularity of the PTP vs. the PWP pattern has been close over the years, with only 2023 being an outlier in terms of PTP pattern dominance. Furthermore, as we shall see, several concrete workflow patterns and methodological choices are furthermore exclusive to either generic workflow pattern. This bisection is therefore in a sense a ”clean cut” and useful for wrapping our heads around an otherwise untamed territory of methodologies.
Figure 12: Line chart illustrating the relative popularity of the PTP vs. the PWP pattern over time in terms of the number of published papers using each pattern.
As the reader may be aware, the notion of a Predict-Then-Optimize workflow pattern already exists within the methodological PSA literature [26], with the same meaning as PTP for mathematical optimization specifically. PTP is simply a generalization to prescriptive method types beyond OPT, according to the diversity of prescriptive methods found in this survey (cf. Section 9). The PWP pattern has to our knowledge never been explicitly named and subjected to formal analysis - possibly since such workflows form newer and rarer contributions to PSA research.
@@
10.3 Predict-Then-Prescribe (PTP) workflows
61 surveyed papers, or 58.7% of all surveyed papers, use a PTP workflow in their solutions. Accounting for 39 of these papers, the most widespread instance of PTP by far is the well-known Predict-Then-Optimize variant [26], in which the role of predictive models, usually based on DM/ML (cf. Figure 9), is to provide inputs to a subsequent prescriptive OPT phase, using Exhaustive Search [140, 92], Dynamic Programming [96, 122], Linear Programming [150, 59, 121], Mixed-Integer Programming [88, 135, 76, 73, 107, 118, 102], Integer Programming [90, 149, 87, 97, 84, 55, 99, 144, 79, 72, 98], Convex Programming [126, 141, 127, 134, 133], Non-Convex Programming [136, 106], or Heuristic Optimization [110, 108, 103] (cf. Section 9.2). In concrete terms, this means that the role of the predictive model is to fill in initial blank coefficients in the optimization model, representing demand forecasts or otherwise, before an optimization algorithm is utilized.